OpenStudy (anonymous):

NEED HELP for this function : 7x√x+4

OpenStudy (anonymous):

wat u hv to find?

OpenStudy (anonymous):

the derivative!

OpenStudy (anonymous):

Use product rule.

OpenStudy (cwrw238):

first change the radical to a fractional index

OpenStudy (anonymous):

yeah i did that but i dont understand it doesnt give me the right answer according to my textbook

OpenStudy (anonymous):

\[ (7x\cdot \sqrt{x+4})'=(7x)'\cdot \sqrt{x+4}+7x\cdot (\sqrt{x+4})' \]

OpenStudy (anonymous):

yeah exactly

OpenStudy (cwrw238):

oh i misread the question - it would have been better if u had writn it as 7x sqrt(x + 4)

OpenStudy (anonymous):

my answer was 49/\[7\sqrt{x+h} +7\sqrt{x}\]

OpenStudy (anonymous):

but in my textbook it says that 49 isnt good , i dont understand